# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 3.1 Auxiliary results
3.2 Analysis of the micro model
@@
where $A\subseteq\partial\Omega^{f}_{\varepsilon}$. To shorten the notation we use, when applicable, the usual notation $H^{1}_{0}(\Omega^{f}_{\varepsilon})=H^{1}_{\partial\Omega^{f}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$. For the solutions of the temperature space we define $\Theta_{\varepsilon}\coloneqq H^{1}(\Omega^{s}_{\varepsilon})\times H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$. For the fluid velocity and pressure we consider
$V_{\varepsilon}\coloneqq\left\{v\in L^{\infty}(S;H^{1}(\Omega^{f}_{\varepsilon})^{d}):v=u_{\varepsilon,BC}\text{ on }\partial\Omega^{f}_{\varepsilon}\text{ and }\operatorname{div}{v}=0\text{ a.e. in }\Omega^{f}_{\varepsilon}\right\}\quad\text{ and }\quad Q_{\varepsilon}\coloneqq L^{\infty}(S;L^{2}_{0}(\Omega^{f}_{\varepsilon})).$
The weak formulation of the heat equation, for a fixed $v\in V_{\varepsilon}$, is given by: Find $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})$ such that $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ and for all $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$
$\begin{split}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}\\ &+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}=(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}}.\end{split}$ (3.1)
@@
$\begin{split}&(\mu(\psi_{\varepsilon})\nabla u_{\varepsilon},\nabla v)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{(v)})_{\Omega^{f}_{\varepsilon}}=0.\end{split}$ (3.2)
For a fixed $\varepsilon>0$ we now prove that there exists a solution of the coupled nonlinear system given by Eq. 3.1 and (3.2). To show this, we follow a similar strategy as in [50], where we split up the problem, first consider both weak formulations on their own, and then show a solution of the coupled system by a fix-point argument of an iterative scheme.
@@
Let the Assumptions (A1), (A3) and (A4) be fulfilled. For a fixed $\varepsilon>0$ and a given $v\in V_{\varepsilon}$ there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ satisfying $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$, $\theta_{\varepsilon}(0,\cdot)=\theta_{\varepsilon,0}$ a.e. and Eq. 3.1. In addition, it holds
$\begin{split}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{s}_{\varepsilon}))}+\frac{1}{\sqrt{\varepsilon}}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{s}_{\varepsilon}))}+\frac{1}{\sqrt{\varepsilon}}\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(S\times\Gamma_{\varepsilon})}\leq C,\end{split}$ (3.3)
@@
We start by showing the existence and uniqueness of a solution, for a fixed $v\in V_{\varepsilon}$. First, note that our problem is of the form
$\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+A_{\varepsilon}(\theta_{\varepsilon},\varphi;t)=(f_{\varepsilon},\varphi^{s})_{\Gamma_{\varepsilon}},$ (3.4)
@@
for a $C<\infty$. Additionally, via integration by parts we have (suppressing the time variable $t$)
$(v\cdot\nabla\psi,\varphi)_{\Omega^{f}_{\varepsilon}}=-(v\cdot\nabla\varphi,\psi)_{\Omega^{f}_{\varepsilon}}-(\operatorname{div}{(v)}\psi,\varphi)_{\Omega^{f}_{\varepsilon}}+(v\cdot n_{\varepsilon},\psi\varphi)_{\partial\Omega^{f}_{\varepsilon}}.$ (3.5)
Plugging in $\varphi=\psi$, that $\operatorname{div}{(v)}=0$ and the boundary conditions on $\partial\Omega^{f}_{\varepsilon}$ into Eq. 3.5 leads to
$(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}=\frac{1}{2}\int_{\Sigma^{\text{out}}_{\varepsilon}}u_{\varepsilon,BC}\cdot n_{\varepsilon}\,\psi^{2}\,\mathrm{d}\sigma_{x}\geq 0$
@@
for all $\psi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$. Given the representation (3.4) and the above estimates for the bilinear form $A_{\varepsilon}$, there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ of the problem [52, Chapter 3.2].
For the norm estimate, we utilize the above computations of the convection term, use $\varphi=\theta_{\varepsilon}$ in the weak formulation and integrate afterward over the time interval $(0,t)$, for $t>0$, to obtain
$\begin{split}\frac{1}{2}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\\ &\leq C\|f_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{s}_{\varepsilon}))}+\frac{1}{2}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}.\end{split}$
Applying Young’s inequality followed by Gronwall’s inequality leads to the desired estimate (3.3), where $C$ is independent of $\varepsilon$. ∎
Remark 3.
Note, the condition in (2.3) is used in the previous Lemma to derive that $(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}\geq 0$, this is not a necessary condition to derive coercivity and carry out the energy estimates. Given the trace inequality and the compact embedding of $H^{1}(\Omega^{f}_{\varepsilon})$ into $L^{2}(\partial\Omega^{f}_{\varepsilon})$, the terms stemming from the convection terms (after integration by part) could be included in the already present gradient terms and $H^{1}$-norm. But (2.3) is a reasonable assumption in our case and let us simplify the notation.
@@
Let the same assumptions be fulfilled as in Lemma 4 and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq. 3.1. Then there exists a $C>0$ independent of $\varepsilon$ such that $\|\theta_{\varepsilon}\|_{L^{\infty}(S\times\Omega)}\leq C.$
@@
For any $k>0$ construct the functions
@@
Note that it holds $\partial_{t}\theta_{\varepsilon,k}^{s}=\chi_{A_{\varepsilon,k}^{s}}\partial_{t}\theta_{\varepsilon}^{s}$ and $\nabla\theta_{\varepsilon,k}^{s}=\chi_{A_{\varepsilon,k}^{s}}\nabla\theta_{\varepsilon}^{s}$ (similiar for $\theta_{\varepsilon,k}^{f}$). By Assumption (A3) we set $k>C_{0}=\sup_{\varepsilon>0}\|\theta_{\varepsilon,0}\|_{L^{\infty}(\Omega)}$. Next, we take $\theta_{\varepsilon,k}=(\theta_{\varepsilon,k}^{s},\theta_{\varepsilon,k}^{f})$ as a test function in Eq. 3.1. Integrating over the time interval $(0,t)\subset S$, we obtain:
$\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(f_{\varepsilon},\theta_{\varepsilon,k}^{s})_{(0,t)\times\Gamma_{\varepsilon}}$
Since $\nabla\theta_{\varepsilon,k}^{f}=\chi_{A_{\varepsilon,k}^{f}}\nabla\theta_{\varepsilon}^{f}$, the convection is
@@
Eq. 3.5 and $\operatorname{div}{v}=0$ yield
$\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot\nu\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot\nu\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$

### v3

@@ 3.1 Auxiliary results
To show the $L^{\infty}$ estimate of the temperature field, we use a truncation argument. In the following lemma, we state some properties of the time derivative of the truncated function in the case that the original time derivative is only an element in the dual space. We denote by $\langle\cdot,\cdot\rangle_{X^{\prime}}$ the dual pairing between elements from the space $X^{\prime}$ and $X$.
Lemma 4.
Let $\psi\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$. For $k>0$ define the truncated function
$\psi_{k}(t,x)\coloneqq\begin{cases}\psi(t,x)-k&\text{if }\psi(t,x)>k,\\ 0&\text{otherwise}.\end{cases}$
Then $\psi_{k}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$. Furthermore, for all $\phi\in L^{2}(S,H^{1}(\Omega))$ with $\phi(t,x)=0$ if $\psi(t,x)\geq k$, it holds $\langle\partial_{t}\psi_{k},\phi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}=0$. If additionally $\psi(0)\in L^{\infty}(\Omega)$. Then for $k\geq\|\psi(0)\|_{L^{\infty}(\Omega)}$ and almost all $t\in S$ it holds
$\int_{0}^{t}\langle\partial_{t}\psi(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau=\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau=\frac{1}{2}(\psi_{k}(t),\psi_{k}(t))_{L^{2}(\Omega)}.$ (4)
Proof.
That $\psi_{k}\in L^{2}(S;H^{1}(\Omega))\cap H^{1}(S;H^{1}(\Omega)^{\prime})$ follows via an approximation with a smooth mollification of $\psi$ and that the cut-off function is Lipschitz continuous. This mollification also directly yields the property $\langle\partial_{t}\psi_{k},\phi\rangle_{L^{2}(S;H^{1}(\Omega)^{\prime})}=0$ for $\phi=0$ if $\psi\geq k$.
For the integral equality (4), we use the fundamental theorem of calculus (with $k\geq\|\psi(0)\|_{L^{\infty}(\Omega)}$) to obtain
$(\psi_{k}(t),\psi_{k}(t))_{L^{2}(\Omega)}=2\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau.$ (5)
Furthermore, it holds
$\begin{split}(\psi_{k}(t),\psi_{k}(t))_{L^{2}(\Omega)}&=(\chi_{\{\psi(t)>k\}}(\psi(t)-k),\psi_{k}(t))_{L^{2}(\Omega)}\\ &=(\psi(t)-k,\psi_{k}(t))_{L^{2}(\Omega)}\\ &=\int_{0}^{t}\langle\partial_{t}\psi(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau+\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi(\tau)-k\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau.\end{split}$ (6)
In the last integral, we can split $\psi-k$ into positive and negative parts. Since $(\psi-k)_{-}=0$ if $\psi\geq k$, this leads to
$\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi(\tau)-k\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau$ $=\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),(\psi(\tau)-k)_{+}\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau+\underbrace{\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),(\psi(\tau)-k)_{-}\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau}_{=0}$
$=\int_{0}^{t}\langle\partial_{t}\psi_{k}(\tau),\psi_{k}(\tau)\rangle_{H^{1}(\Omega)^{\prime}}\,\mathrm{d}\tau.$
Plugging the last equation into Equation 6 and comparing with Equation 5 yields the desired result. ∎
3.2 Well-posedness and a prior estimates
@@
where $A\subseteq\partial\Omega^{f}_{\varepsilon}$. To shorten the notation we use, when applicable, the usual notation $H^{1}_{0}(\Omega^{f}_{\varepsilon})=H^{1}_{\partial\Omega^{f}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$. For the solutions of the temperature space we define $\Theta_{\varepsilon}\coloneqq H^{1}(\Omega^{s}_{\varepsilon})\times H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})$. The dual of a finite product space is isometrically isomorph to the product of the dual spaces. For any element $\psi\in\Theta_{\varepsilon}^{\prime}$, there are unique $\psi^{s}\in H^{1}(\Omega^{s}_{\varepsilon})^{\prime}$, $\psi^{f}\in H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}$ such that for all $\varphi\in\Theta_{\varepsilon}$
$\langle\psi,\varphi\rangle_{\Theta_{\varepsilon}^{\prime}}=\langle\psi^{s},\varphi^{s}\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\langle\psi^{f},\varphi^{f}\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}.$
In the following, we identify $\psi=(\psi^{s},\psi^{f})\in\Theta_{\varepsilon}^{\prime}$ and suppress the superscript when the context is clear from the dual pairing.
For the fluid velocity and pressure, we define
@@
$\hskip 85.35826ptQ_{\varepsilon}\coloneqq L^{\infty}(S;L^{2}_{0}(\Omega^{f}_{\varepsilon})).$
The weak formulation of the heat equation, for a fixed $v\in V_{\varepsilon}$, is given by: Find $\theta_{\varepsilon}=(\theta^{s}_{\varepsilon},\theta^{f}_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})$ such that $\partial_{t}\theta_{\varepsilon}=(\partial_{t}\theta_{\varepsilon}^{s},\partial_{t}\theta_{\varepsilon}^{f})\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ and for all $\varphi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$
$\begin{split}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{s}_{\varepsilon}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}(\kappa\nabla\theta_{\varepsilon},\nabla\varphi)_{\Omega^{f}_{\varepsilon}}\\ &+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\varphi)_{\Omega^{f}_{\varepsilon}}+\alpha(\llbracket\theta_{\varepsilon}\rrbracket,\llbracket\varphi\rrbracket)_{\Gamma_{\varepsilon}}=(f_{\varepsilon}^{s},\varphi^{s})_{\Gamma_{\varepsilon}}+(f_{\varepsilon}^{f},\varphi^{f})_{\Gamma_{\varepsilon}}.\end{split}$ (7)
Here, $(\cdot,\cdot)_{D}$ denotes the usual $L^{2}$ inner product over the domain $D$. Given a $\psi_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$, the weak form of the Stokes equation is: Find $(u_{\varepsilon},p_{\varepsilon})\in V_{\varepsilon}\times Q_{\varepsilon}$ such that for all $v\in H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$ and almost all $t\in S$
$\begin{split}&(\mu(\psi_{\varepsilon})\nabla u_{\varepsilon},\nabla v)_{\Omega^{f}_{\varepsilon}}-(p_{\varepsilon},\operatorname{div}{(v)})_{\Omega^{f}_{\varepsilon}}=0.\end{split}$ (8)
For a fixed $\varepsilon>0$ we now prove that there exists a solution of the coupled nonlinear system given by Eq. 7 and (8). To show this, we follow a similar strategy as in [6], where we split up the problem, first consider both weak formulations on their own, and then show a solution of the coupled system by a fix-point argument of an iterative scheme.
@@
Let the Assumptions (A1), (A3) and (A4) be fulfilled. For a fixed $\varepsilon>0$ and a given $v\in V_{\varepsilon}$ there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ satisfying $\partial_{t}\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$, $\theta_{\varepsilon}(0,\cdot)=\theta_{\varepsilon,0}$ a.e. and Eq. 7. In addition, it holds
$\begin{split}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{s}_{\varepsilon}))}&+\frac{1}{\sqrt{\varepsilon}}\|\theta_{\varepsilon}\|_{L^{\infty}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{s}_{\varepsilon}))}\\ &+\frac{1}{\sqrt{\varepsilon}}\|\nabla\theta_{\varepsilon}\|_{L^{2}(S;L^{2}(\Omega^{f}_{\varepsilon}))}+\|\llbracket\theta_{\varepsilon}\rrbracket\|_{L^{2}(S\times\Gamma_{\varepsilon})}\leq C,\end{split}$ (9)
@@
We start by showing the existence and uniqueness of a solution for a fixed $v\in V_{\varepsilon}$. First, note that our problem is of the form
$\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}+\frac{1}{\varepsilon}\langle\partial_{t}\theta_{\varepsilon},\varphi\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+A_{\varepsilon}(\theta_{\varepsilon},\varphi;t)=(f_{\varepsilon}^{s},\varphi^{s})_{\Gamma_{\varepsilon}}+(f_{\varepsilon}^{f},\varphi^{f})_{\Gamma_{\varepsilon}},$ (10)
@@
for a $C<\infty$. Additionally, via integration by parts, we have (suppressing the time variable $t$)
$(v\cdot\nabla\psi,\varphi)_{\Omega^{f}_{\varepsilon}}=-(v\cdot\nabla\varphi,\psi)_{\Omega^{f}_{\varepsilon}}-(\operatorname{div}{(v)}\psi,\varphi)_{\Omega^{f}_{\varepsilon}}+(v\cdot n,\psi\varphi)_{\partial\Omega^{f}_{\varepsilon}}.$ (11)
Plugging in $\varphi=\psi$, that $\operatorname{div}{(v)}=0$ and the boundary conditions on $\partial\Omega^{f}_{\varepsilon}$ into Eq. 11 leads to
$(v\cdot\nabla\psi,\psi)_{\Omega^{f}_{\varepsilon}}=\frac{1}{2}\int_{\Sigma^{\text{out}}_{\varepsilon}}u_{\varepsilon,BC}\cdot n\,\psi^{2}\,\mathrm{d}\sigma_{x}\geq 0$
@@
for all $\psi\in L^{2}(S;\Theta_{\varepsilon})$ and almost all $t\in S$. Given the representation (10) and the above estimates for the bilinear form $A_{\varepsilon}$, there exists a unique solution $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ [46, Chapter 3, Proposition 2.3].
For the norm estimate, we utilize the above computations of the convection term, use $\varphi=\theta_{\varepsilon}$ in the weak formulation, and integrate afterward over the time interval $(0,t)$ to obtain (with the trace estimates from Lemma 2)
$\begin{split}\frac{1}{2}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}&+\frac{1}{2\varepsilon}\|\theta_{\varepsilon}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon}\|_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}^{2}+\alpha\|\llbracket\theta_{\varepsilon}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\\ &\leq C\|f^{s}_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{s}_{\varepsilon}))}+\frac{C}{\sqrt{\varepsilon}}\|f^{f}_{\varepsilon}\|_{L^{2}((0,t)\times\Gamma_{\varepsilon})}\|\theta_{\varepsilon}\|_{L^{2}((0,t);H^{1}(\Omega^{f}_{\varepsilon}))}\\ &\hskip 10.0pt+\frac{1}{2}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,0}\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}.\end{split}$
Applying Young’s inequality followed by Gronwall’s inequality leads to the desired estimate (9), where $C$ is independent of $\varepsilon$. ∎
@@
Let the same assumptions as in Lemma 5 be fulfilled and let $\theta_{\varepsilon}\in L^{2}(S;\Theta_{\varepsilon})$ be the solution of Eq. 7. Then there exists a $C>0$ independent of $\varepsilon$ such that $\|\theta_{\varepsilon}\|_{L^{\infty}(S\times\Omega)}\leq C.$
@@
To show the estimate, we observe that the solution is linear in the source terms $f_{\varepsilon}^{s,f}$. Therefore, we can show the estimates for the two cases $f_{\varepsilon}^{s}=0,f_{\varepsilon}^{f}\neq 0$ and $f_{\varepsilon}^{s}\neq 0,f_{\varepsilon}^{f}=0$ separately. Both cases follow via similar arguments, where we present the case $f_{\varepsilon}^{s}=0,f_{\varepsilon}^{f}\neq 0$ and in the end shortly comment on the other case. To avoid cluttering, we do not introduce an additional index to denote this case.
First, for any $k>0$ construct the functions
@@
Note $\nabla\theta_{\varepsilon,k}^{s,f}=\chi_{A_{\varepsilon,k}^{s,f}}\nabla\theta_{\varepsilon,k}^{s,f}$. Since $\theta_{\varepsilon,k}^{s,f}=(\theta_{\varepsilon}^{s,f}-k)_{+}+(\theta_{\varepsilon}^{s,f}+k)_{-}$, applying Lemma 4 to both the positive and negative part yields, $\partial_{t}\theta_{\varepsilon,k}^{s,f}\in L^{2}(S;H^{1}(\Omega^{s,f}_{\varepsilon})^{\prime})$. By Assumption (A3) we set $k>C_{0}=\sup_{\varepsilon>0}\|\theta_{\varepsilon,0}\|_{L^{\infty}(\Omega)}$. Next, we take $\theta_{\varepsilon,k}=(\theta_{\varepsilon,k}^{s},\theta_{\varepsilon,k}^{f})$ as a test function in Eq. 7, leading, for almost all $\tau\in S$, to
$\begin{split}\langle\partial_{t}\theta^{s}_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{s}(\tau)\rangle_{H^{1}(\Omega^{s}_{\varepsilon})^{\prime}}&+\frac{1}{\varepsilon}\langle\partial_{t}\theta^{f}_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{f}(\tau)\rangle_{H^{1}_{\Sigma^{\text{in}}_{\varepsilon}}(\Omega^{f}_{\varepsilon})^{\prime}}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}(\tau)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}(\tau)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}\\ &+\frac{1}{\varepsilon}(v(\tau)\cdot\nabla\theta_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{f}(\tau))_{\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}(\tau)\rrbracket\|^{2}_{L^{2}(\Gamma_{\varepsilon})}=(f^{f}_{\varepsilon}(\tau),\theta_{\varepsilon,k}^{f}(\tau))_{\Gamma_{\varepsilon}}.\end{split}$
Integrating over the time interval $(0,t)\subset S$ and using Equation 4 from Lemma 4 we obtain
$\frac{1}{2}\|\theta_{\varepsilon,k}^{s}(t)\|^{2}_{L^{2}(\Omega^{s}_{\varepsilon})}$ $+\frac{1}{2\varepsilon}\|\theta_{\varepsilon,k}^{f}(t)\|^{2}_{L^{2}(\Omega^{f}_{\varepsilon})}+\kappa^{s}\|\nabla\theta_{\varepsilon,k}^{s}\|^{2}_{L^{2}((0,t)\times\Omega^{s}_{\varepsilon})}+\frac{1}{\varepsilon}\kappa^{f}\|\nabla\theta_{\varepsilon,k}^{f}\|^{2}_{L^{2}((0,t)\times\Omega^{f}_{\varepsilon})}$
$+\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}+\alpha\|\llbracket\theta_{\varepsilon,k}\rrbracket\|^{2}_{L^{2}((0,t)\times\Gamma_{\varepsilon})}=(f_{\varepsilon}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Gamma_{\varepsilon}}$
Since $\nabla\theta_{\varepsilon,k}^{f}=\chi_{A_{\varepsilon,k}^{f}}\nabla\theta_{\varepsilon}^{f}$, for the the convection it holds
@@
Eq. 11 and $\operatorname{div}{v}=0$ yield
$\frac{1}{\varepsilon}(v\cdot\nabla\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot n\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Omega^{f}_{\varepsilon}}=\frac{1}{2\varepsilon}(v\cdot n\theta_{\varepsilon,k}^{f},\theta_{\varepsilon,k}^{f})_{(0,t)\times\partial\Sigma^{\text{out}}_{\varepsilon}}\geq 0.$